# Towards Building Multilingual Language Model for Medicine

### v2

@@ C.2 Different Metrics on Rationale Evaluation
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ BLEU-1 45.94 48.37 44.71 44.47 29.95 45.91 43.22 InternLM BLEU-1 46.53 48.24 44.89 41.80 27.87 43.42 42.12 LLaMA 2 BLEU-1 46.87 46.62 48.53 44.43 33.05 45.96 44.24 ChatDoctor BLEU-1 47.22 44.66 38.87 44.64 32.19 45.68 42.21 MedAlpaca BLEU-1 47.33 45.72 45.35 43.78 32.80 45.99 43.49 PMC-LLaMA BLEU-1 47.33 45.87 44.52 43.80 31.14 46.30 43.16 Mistral BLEU-1 47.16 48.34 48.80 45.83 34.52 47.55 45.37 InternLM 2 BLEU-1 49.48 51.38 50.64 46.73 32.93 47.94 46.52 MMedLM BLEU-1 47.37 48.68 48.95 45.39 33.24 46.68 45.05 MMedLM 2 BLEU-1 50.02 51.39 54.79 49.04 37.49 50.14 48.81 BLOOMZ BLEU-2 17.93 18.70 17.89 18.29 8.58 17.69 16.51 InternLM BLEU-2 18.76 18.79 18.55 15.89 7.44 15.89 15.89 LLaMA 2 BLEU-2 18.53 16.98 20.42 17.72 9.40 17.31 16.73 ChatDoctor BLEU-2 19.50 16.29 15.36 18.25 9.76 17.87 16.17 MedAlpaca BLEU-2 19.30 16.41 18.48 17.82 9.90 17.91 16.64 PMC-LLaMA BLEU-2 20.05 16.28 17.67 17.99 9.10 17.68 16.46 Mistral BLEU-2 18.57 17.97 20.47 19.02 10.48 18.33 17.47 InternLM 2 BLEU-2 21.20 20.76 22.83 20.14 9.72 19.12 18.96 MMedLM BLEU-2 19.39 18.84 21.63 19.46 10.13 18.99 18.07 MMedLM 2 BLEU-2 21.52 21.46 26.75 23.18 13.42 21.54 21.31 BLOOMZ BLEU-3 8.39 8.97 8.90 9.15 3.66 8.42 7.92 InternLM BLEU-3 9.20 9.05 9.56 7.56 3.09 7.24 7.62 LLaMA 2 BLEU-3 8.89 7.64 10.41 8.78 3.63 7.98 7.89 ChatDoctor BLEU-3 9.62 7.34 7.74 9.17 4.10 8.63 7.77 MedAlpaca BLEU-3 9.62 7.66 9.39 8.98 4.16 8.65 8.08 PMC-LLaMA BLEU-3 10.14 7.27 8.91 8.91 3.88 8.46 7.93 Mistral BLEU-3 8.67 8.42 10.33 9.65 4.41 8.76 8.37 InternLM 2 BLEU-3 10.71 10.54 12.41 10.55 4.12 9.31 9.61 MMedLM BLEU-3 9.55 9.20 11.84 10.24 4.20 9.38 9.07 MMedLM 2 BLEU-3 11.00 11.02 15.05 12.99 6.37 11.22 11.27 BLOOMZ BLEU-4 4.55 4.62 4.87 4.73 1.76 4.58 4.18 InternLM BLEU-4 5.23 4.82 5.51 3.88 1.55 3.90 4.15 LLaMA 2 BLEU-4 4.85 3.67 5.94 4.66 1.66 4.28 4.17 ChatDoctor BLEU-4 5.47 3.65 4.42 4.86 1.96 4.71 4.18 MedAlpaca BLEU-4 5.43 3.88 5.33 4.83 1.95 4.89 4.38 PMC-LLaMA BLEU-4 5.79 3.58 5.04 4.52 1.93 4.66 4.25 Mistral BLEU-4 4.75 4.30 5.74 5.16 2.18 4.88 4.50 InternLM 2 BLEU-4 6.13 5.84 7.41 5.83 2.03 5.17 5.40 MMedLM BLEU-4 5.42 5.02 7.09 5.71 1.88 5.39 5.08 MMedLM 2 BLEU-4 6.41 6.20 9.15 7.70 3.30 6.63 6.56 BLOOMZ BLEU 12.80 13.14 13.21 13.06 5.70 12.53 11.74 InternLM BLEU 13.73 13.46 13.95 11.19 4.97 11.17 11.41 LLaMA 2 BLEU 13.32 11.44 15.28 12.72 5.83 12.10 11.78 ChatDoctor BLEU 14.10 11.09 11.49 13.09 6.31 12.81 11.48 MedAlpaca BLEU 14.05 11.49 13.74 12.99 6.36 12.97 11.94 PMC-LLaMA BLEU 14.74 11.01 13.22 12.76 6.06 12.77 11.76 Mistral BLEU 13.11 12.65 15.16 13.77 6.90 13.25 12.47 InternLM 2 BLEU 15.44 15.28 17.59 14.90 6.48 13.80 13.91 MMedLM BLEU 14.01 13.69 16.72 14.38 6.36 13.80 13.16 MMedLM 2 BLEU 15.92 15.80 20.75 17.72 9.35 16.11 15.94
@@
[b]
Method Metric English Chinese Japanese French Russian Spanish Avg.
BLOOMZ ROUGE-1 36.81 45.17 38.09 37.49 15.28 39.22 35.34
InternLM ROUGE-1 38.05 44.71 37.57 32.14 16.79 34.56 33.97
LLaMA 2 ROUGE-1 39.34 43.71 49.53 39.29 22.66 39.64 39.03
ChatDoctor ROUGE-1 40.26 42.80 45.00 39.06 22.84 40.23 38.37
MedAlpaca ROUGE-1 39.52 42.50 44.69 39.57 22.83 39.64 38.12
PMC-LLaMA ROUGE-1 40.90 42.95 43.67 38.64 22.45 39.80 38.07
Mistral ROUGE-1 42.80 46.31 48.19 41.07 24.75 42.83 40.99
InternLM 2 ROUGE-1 41.66 49.28 46.91 38.46 21.71 40.13 39.69
MMedLM ROUGE-1 38.83 46.38 46.90 37.78 21.28 38.77 38.32
MMedLM 2 ROUGE-1 41.51 48.36 52.24 41.36 25.70 42.64 41.97
BLOOMZ ROUGE-2 12.73 16.77 13.43 14.01 4.03 15.66 12.77
InternLM ROUGE-2 13.08 16.71 13.92 10.80 4.43 12.23 11.86
LLaMA 2 ROUGE-2 13.63 15.18 21.37 15.06 6.41 15.13 14.47
ChatDoctor ROUGE-2 15.25 15.10 19.18 15.64 6.82 16.13 14.69
MedAlpaca ROUGE-2 14.77 15.09 18.86 15.68 6.65 15.83 14.48
PMC-LLaMA ROUGE-2 15.59 15.37 18.11 14.81 6.66 15.90 14.41
Mistral ROUGE-2 17.28 17.60 20.49 17.00 7.72 18.10 16.36
InternLM 2 ROUGE-2 15.95 20.04 20.26 14.55 6.00 15.62 15.40
MMedLM ROUGE-2 13.78 17.68 19.75 14.44 6.22 15.27 14.52
MMedLM 2 ROUGE-2 16.02 19.94 23.27 17.01 8.29 18.22 17.12
BLOOMZ ROUGE-L 33.28 30.71 26.91 34.01 14.56 35.22 29.11
InternLM ROUGE-L 34.26 31.02 26.31 29.37 15.76 30.99 27.95
LLaMA 2 ROUGE-L 34.91 28.40 31.82 35.38 21.02 34.99 31.08
ChatDoctor ROUGE-L 36.57 28.64 30.86 35.63 21.15 35.98 31.47
MedAlpaca ROUGE-L 35.88 28.87 30.35 36.08 21.47 35.41 31.34
PMC-LLaMA ROUGE-L 36.84 28.90 29.43 34.96 20.92 35.35 31.07
Mistral ROUGE-L 38.57 31.23 31.49 37.23 23.18 37.90 33.27
InternLM 2 ROUGE-L 37.45 34.57 31.70 34.95 20.23 35.78 32.45
MMedLM ROUGE-L 35.08 32.18 32.13 34.19 19.70 34.41 31.28
MMedLM 2 ROUGE-L 37.61 33.52 33.92 37.32 23.95 37.95 34.04
@@
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ ROUGE-1 40.51 48.26 48.61 41.05 21.50 40.77 40.12 InternLM ROUGE-1 41.86 48.64 49.83 37.95 21.20 38.59 39.68 LLaMA 2 ROUGE-1 41.39 46.57 51.21 40.38 23.24 40.37 40.53 ChatDoctor ROUGE-1 41.97 45.81 47.95 40.25 23.37 40.71 40.01 MedAlpaca ROUGE-1 42.31 46.49 49.12 40.41 23.15 40.57 40.34 PMC-LLaMA ROUGE-1 42.87 46.18 48.44 40.23 22.28 40.68 40.12 Mistral ROUGE-1 41.82 47.91 50.60 40.88 24.68 41.41 41.22 InternLM 2 ROUGE-1 44.12 51.58 53.46 42.00 24.05 41.96 42.86 MMedLM ROUGE-1 41.98 49.28 52.34 41.41 24.67 41.35 41.84 MMedLM 2 ROUGE-1 44.77 51.78 57.10 45.30 28.18 44.59 45.29 BLOOMZ ROUGE-2 15.25 19.78 21.22 16.63 6.13 16.44 15.91 InternLM ROUGE-2 16.19 19.86 22.13 14.67 5.64 14.92 15.57 LLaMA 2 ROUGE-2 15.75 18.07 22.81 15.76 6.33 15.81 15.76 ChatDoctor ROUGE-2 16.70 17.67 20.65 16.16 7.12 16.64 15.82 MedAlpaca ROUGE-2 16.41 17.63 21.78 16.21 7.11 16.49 15.94 PMC-LLaMA ROUGE-2 17.40 17.56 21.00 16.20 6.77 16.47 15.90 Mistral ROUGE-2 15.64 18.71 22.64 16.64 7.31 16.71 16.27 InternLM 2 ROUGE-2 18.04 21.75 25.32 17.76 7.02 17.36 17.88 MMedLM ROUGE-2 16.31 20.00 24.50 17.82 7.57 17.56 17.29 MMedLM 2 ROUGE-2 18.47 22.56 28.88 20.80 10.02 19.68 20.07 BLOOMZ ROUGE-L 36.53 32.84 32.03 37.17 19.90 35.94 32.40 InternLM ROUGE-L 37.89 33.59 32.49 34.37 20.13 34.26 32.12 LLaMA 2 ROUGE-L 37.05 31.09 32.45 36.44 21.73 35.72 32.41 ChatDoctor ROUGE-L 37.91 31.04 31.62 36.29 22.02 36.38 32.54 MedAlpaca ROUGE-L 38.00 31.43 32.08 36.57 21.77 36.13 32.66 PMC-LLaMA ROUGE-L 38.77 30.85 31.42 36.57 21.13 36.19 32.49 Mistral ROUGE-L 37.22 32.06 31.78 36.88 23.00 36.35 32.88 InternLM 2 ROUGE-L 39.61 36.27 35.06 38.17 22.74 37.24 34.85 MMedLM ROUGE-L 37.84 34.35 34.51 37.81 23.19 37.08 34.13 MMedLM 2 ROUGE-L 40.19 36.97 38.31 41.17 26.62 39.71 37.16
@@
[b] Method English Chinese Japanese French Russian Spanish Avg. BLOOMZ 74.72 76.46 67.87 76.36 67.24 75.58 73.04 InternLM 75.13 76.95 67.22 72.46 67.87 72.94 72.09 LLaMA 2 76.50 76.38 74.81 77.77 74.30 76.54 76.05 ChatDoctor 76.53 75.23 72.31 77.34 74.25 76.66 75.39 MedAlpaca 76.30 75.43 71.33 77.36 74.14 76.43 75.16 PMC-LLaMA 76.84 75.17 71.28 77.15 73.77 76.35 75.09 Mistral 77.99 77.46 74.11 78.67 75.07 77.77 76.85 InternLM 2 77.36 78.75 73.05 77.58 73.68 76.66 76.18 MMedLM 75.68 77.16 72.15 76.41 73.87 75.42 75.12 MMedLM 2 77.59 78.42 75.79 78.47 76.07 77.65 77.33
@@
[b] Method English Chinese Japanese French Russian Spanish Avg. BLOOMZ 76.66 78.30 74.26 78.18 72.22 76.92 76.09 InternLM 77.11 78.43 74.73 77.04 71.91 76.04 75.88 LLaMA 2 77.49 77.78 75.99 78.45 74.60 77.21 76.92 ChatDoctor 77.47 77.18 72.65 78.16 74.69 77.19 76.22 MedAlpaca 77.53 77.31 74.39 78.23 74.85 77.24 76.59 PMC-LLaMA 77.98 77.22 74.15 78.15 73.94 76.96 76.40 Mistral 77.79 78.49 76.00 79.00 75.77 77.80 77.47 InternLM 2 78.44 79.37 76.70 79.38 74.95 78.01 77.81 MMedLM 77.58 78.66 76.02 78.71 75.42 77.36 77.29 MMedLM 2 78.74 79.56 78.40 80.36 77.41 78.73 78.87
@@
C.3 Detailed Results on Rating Scores
In Table 14 and 15, we further show the GPT-4 and human rating scores for six languages respectively.

### v3

@@ E.2 Different Metrics on Rationale Evaluation
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ BLEU-1 45.94 48.37 44.71 44.47 29.95 45.91 43.22 InternLM BLEU-1 46.53 48.24 44.89 41.80 27.87 43.42 42.12 Llama 2 BLEU-1 46.87 46.62 48.53 44.43 33.05 45.96 44.24 ChatDoctor BLEU-1 47.22 44.66 38.87 44.64 32.19 45.68 42.21 MedAlpaca BLEU-1 47.33 45.72 45.35 43.78 32.80 45.99 43.49 PMC-LLaMA BLEU-1 47.33 45.87 44.52 43.80 31.14 46.30 43.16 Mistral BLEU-1 47.16 48.34 48.80 45.83 34.52 47.55 45.37 Meditron BLEU-1 47.40 47.93 49.13 45.93 33.65 46.42 45.08 InternLM 2 BLEU-1 49.48 51.38 50.64 46.73 32.93 47.94 46.52 BioMistral BLEU-1 47.96 49.76 49.73 46.34 34.20 47.57 45.93 Llama 3 BLEU-1 48.74 49.44 51.96 47.11 34.73 48.59 46.76 MMedLM (Ours) BLEU-1 47.37 48.68 48.95 45.39 33.24 46.68 45.05 MMedLM 2 (Ours) BLEU-1 50.02 51.39 54.79 49.04 37.49 50.14 48.81 MMed-Llama 3 (Ours) BLEU-1 47.61 49.96 52.89 47.92 36.32 48.61 47.22 BLOOMZ BLEU-2 17.93 18.70 17.89 18.29 8.58 17.69 16.51 InternLM BLEU-2 18.76 18.79 18.55 15.89 7.44 15.89 15.89 Llama 2 BLEU-2 18.53 16.98 20.42 17.72 9.40 17.31 16.73 ChatDoctor BLEU-2 19.50 16.29 15.36 18.25 9.76 17.87 16.17 MedAlpaca BLEU-2 19.30 16.41 18.48 17.82 9.90 17.91 16.64 PMC-LLaMA BLEU-2 20.05 16.28 17.67 17.99 9.10 17.68 16.46 Mistral BLEU-2 18.57 17.97 20.47 19.02 10.48 18.33 17.47 Meditron BLEU-2 19.86 18.29 20.92 19.16 10.04 18.43 17.78 InternLM 2 BLEU-2 21.20 20.76 22.83 20.14 9.72 19.12 18.96 BioMistral BLEU-2 18.91 19.59 21.73 19.65 10.45 18.54 18.15 Llama 3 BLEU-2 20.80 19.29 23.48 20.50 10.78 20.26 19.19 MMedLM (Ours) BLEU-2 19.39 18.84 21.63 19.46 10.13 18.99 18.07 MMedLM 2 (Ours) BLEU-2 21.52 21.46 26.75 23.18 13.42 21.54 21.31 MMed-Llama 3 (Ours) BLEU-2 20.07 19.78 24.73 20.95 11.92 20.31 19.63 BLOOMZ BLEU-3 8.39 8.97 8.90 9.15 3.66 8.42 7.92 InternLM BLEU-3 9.20 9.05 9.56 7.56 3.09 7.24 7.62 Llama 2 BLEU-3 8.89 7.64 10.41 8.78 3.63 7.98 7.89 ChatDoctor BLEU-3 9.62 7.34 7.74 9.17 4.10 8.63 7.77 MedAlpaca BLEU-3 9.62 7.66 9.39 8.98 4.16 8.65 8.08 PMC-LLaMA BLEU-3 10.14 7.27 8.91 8.91 3.88 8.46 7.93 Mistral BLEU-3 8.67 8.42 10.33 9.65 4.41 8.76 8.37 Meditron BLEU-3 10.02 8.75 10.81 9.79 4.19 8.98 8.76 InternLM 2 BLEU-3 10.71 10.54 12.41 10.55 4.12 9.31 9.61 BioMistral BLEU-3 8.96 9.73 11.59 10.14 4.76 9.07 9.04 Llama 3 BLEU-3 10.69 9.32 12.65 10.98 4.54 10.56 9.79 MMedLM (Ours) BLEU-3 9.55 9.20 11.84 10.24 4.20 9.38 9.07 MMedLM 2 (Ours) BLEU-3 11.00 11.02 15.05 12.99 6.37 11.22 11.27 MMed-Llama 3 (Ours) BLEU-3 10.21 9.90 13.41 11.11 5.37 10.58 10.10 BLOOMZ BLEU-4 4.55 4.62 4.87 4.73 1.76 4.58 4.18 InternLM BLEU-4 5.23 4.82 5.51 3.88 1.55 3.90 4.15 Llama 2 BLEU-4 4.85 3.67 5.94 4.66 1.66 4.28 4.17 ChatDoctor BLEU-4 5.47 3.65 4.42 4.86 1.96 4.71 4.18 MedAlpaca BLEU-4 5.43 3.88 5.33 4.83 1.95 4.89 4.38 PMC-LLaMA BLEU-4 5.79 3.58 5.04 4.52 1.93 4.66 4.25 Meditron BLEU-4 5.77 4.53 6.09 5.44 2.04 4.99 4.81 Mistral BLEU-4 4.75 4.30 5.74 5.16 2.18 4.88 4.50 BioMistral BLEU-4 4.96 5.33 6.76 5.57 2.51 5.05 5.03 Llama 3 BLEU-4 6.41 4.80 7.31 6.19 2.17 6.27 5.53 InternLM 2 BLEU-4 6.13 5.84 7.41 5.83 2.03 5.17 5.40 MMedLM (Ours) BLEU-4 5.42 5.02 7.09 5.71 1.88 5.39 5.08 MMedLM 2 (Ours) BLEU-4 6.41 6.20 9.15 7.70 3.30 6.63 6.56 MMed-Llama 3 (Ours) BLEU-4 5.92 5.43 7.79 6.30 2.64 6.17 5.71 BLOOMZ BLEU 12.80 13.14 13.21 13.06 5.70 12.53 11.74 InternLM BLEU 13.73 13.46 13.95 11.19 4.97 11.17 11.41 Llama 2 BLEU 13.32 11.44 15.28 12.72 5.83 12.10 11.78 ChatDoctor BLEU 14.10 11.09 11.49 13.09 6.31 12.81 11.48 MedAlpaca BLEU 14.05 11.49 13.74 12.99 6.36 12.97 11.94 PMC-LLaMA BLEU 14.74 11.01 13.22 12.76 6.06 12.77 11.76 Meditron BLEU 14.61 13.02 15.68 14.02 6.57 13.28 12.86 Mistral BLEU 13.11 12.65 15.16 13.77 6.90 13.25 12.47 BioMistral BLEU 13.54 14.23 16.55 14.37 7.23 13.52 13.24 Llama 3 BLEU 15.61 13.67 17.87 15.29 6.92 15.27 14.11 InternLM 2 BLEU 15.44 15.28 17.59 14.90 6.48 13.80 13.91 MMedLM (Ours) BLEU 14.01 13.69 16.72 14.38 6.36 13.80 13.16 MMedLM 2 (Ours) BLEU 15.92 15.80 20.75 17.72 9.35 16.11 15.94 MMed-Llama 3 (Ours) BLEU 14.87 14.39 18.75 15.60 7.96 15.21 14.46
@@
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ ROUGE-1 36.81 45.17 38.09 37.49 15.28 39.22 35.34 InternLM ROUGE-1 38.05 44.71 37.57 32.14 16.79 34.56 33.97 Llama 2 ROUGE-1 39.34 43.71 49.53 39.29 22.66 39.64 39.03 ChatDoctor ROUGE-1 40.26 42.80 45.00 39.06 22.84 40.23 38.37 MedAlpaca ROUGE-1 39.52 42.50 44.69 39.57 22.83 39.64 38.12 PMC-LLaMA ROUGE-1 40.90 42.95 43.67 38.64 22.45 39.80 38.07 Meditron ROUGE-1 40.42 42.06 43.34 37.88 21.64 39.06 37.40 Mistral ROUGE-1 42.80 46.31 48.19 41.07 24.75 42.83 40.99 BioMistral ROUGE-1 41.34 44.75 44.46 39.41 18.80 41.07 38.31 InternLM 2 ROUGE-1 41.66 49.28 46.91 38.46 21.71 40.13 39.69 Llama 3 ROUGE-1 41.73 47.44 50.43 40.69 22.22 42.70 40.87 MMedLM (Ours) ROUGE-1 38.83 46.38 46.90 37.78 21.28 38.77 38.32 MMedLM 2 (Ours) ROUGE-1 41.51 48.36 52.24 41.36 25.70 42.64 41.97 MMed-Llama 3 (Ours) ROUGE-1 41.57 47.71 53.18 40.97 24.87 43.32 41.94 BLOOMZ ROUGE-2 12.73 16.77 13.43 14.01 4.03 15.66 12.77 InternLM ROUGE-2 13.08 16.71 13.92 10.80 4.43 12.23 11.86 Llama 2 ROUGE-2 13.63 15.18 21.37 15.06 6.41 15.13 14.47 ChatDoctor ROUGE-2 15.25 15.10 19.18 15.64 6.82 16.13 14.69 MedAlpaca ROUGE-2 14.77 15.09 18.86 15.68 6.65 15.83 14.48 PMC-LLaMA ROUGE-2 15.59 15.37 18.11 14.81 6.66 15.90 14.41 Meditron ROUGE-2 15.12 15.03 16.28 14.15 6.07 14.93 13.60 Mistral ROUGE-2 17.28 17.60 20.49 17.00 7.72 18.10 16.36 BioMistral ROUGE-2 16.17 16.86 17.96 15.30 4.38 16.57 14.54 InternLM 2 ROUGE-2 15.95 20.04 20.26 14.55 6.00 15.62 15.40 Llama 3 ROUGE-2 16.38 18.79 22.16 16.02 6.32 17.99 16.28 MMedLM (Ours) ROUGE-2 13.78 17.68 19.75 14.44 6.22 15.27 14.52 MMedLM 2 (Ours) ROUGE-2 16.02 19.94 23.27 17.01 8.29 18.22 17.12 MMed-Llama 3 (Ours) ROUGE-2 16.26 18.81 24.53 17.06 7.87 18.92 17.24 BLOOMZ ROUGE-L 33.28 30.71 26.91 34.01 14.56 35.22 29.11 InternLM ROUGE-L 34.26 31.02 26.31 29.37 15.76 30.99 27.95 Llama 2 ROUGE-L 34.91 28.40 31.82 35.38 21.02 34.99 31.08 ChatDoctor ROUGE-L 36.57 28.64 30.86 35.63 21.15 35.98 31.47 MedAlpaca ROUGE-L 35.88 28.87 30.35 36.08 21.47 35.41 31.34 PMC-LLaMA ROUGE-L 36.84 28.90 29.43 34.96 20.92 35.35 31.07 Meditron ROUGE-L 36.38 28.58 29.25 34.24 20.19 34.61 30.54 Mistral ROUGE-L 38.57 31.23 31.49 37.23 23.18 37.90 33.27 BioMistral ROUGE-L 37.13 30.15 29.79 35.58 17.57 36.36 31.10 InternLM 2 ROUGE-L 37.45 34.57 31.70 34.95 20.23 35.78 32.45 Llama 3 ROUGE-L 37.91 32.71 33.36 36.55 20.79 37.86 33.19 MMedLM (Ours) ROUGE-L 35.08 32.18 32.13 34.19 19.70 34.41 31.28 MMedLM 2 (Ours) ROUGE-L 37.61 33.52 33.92 37.32 23.95 37.95 34.04 MMed-Llama 3 (Ours) ROUGE-L 37.26 33.05 35.02 37.57 23.37 38.49 34.13
@@
[b] Method Metric English Chinese Japanese French Russian Spanish Avg. BLOOMZ ROUGE-1 40.51 48.26 48.61 41.05 21.50 40.77 40.12 InternLM ROUGE-1 41.86 48.64 49.83 37.95 21.20 38.59 39.68 Llama 2 ROUGE-1 41.39 46.57 51.21 40.38 23.24 40.37 40.53 ChatDoctor ROUGE-1 41.97 45.81 47.95 40.25 23.37 40.71 40.01 MedAlpaca ROUGE-1 42.31 46.49 49.12 40.41 23.15 40.57 40.34 PMC-LLaMA ROUGE-1 42.87 46.18 48.44 40.23 22.28 40.68 40.12 Meditron ROUGE-1 42.85 48.61 52.03 41.37 24.10 41.11 41.68 Mistral ROUGE-1 41.82 47.91 50.60 40.88 24.68 41.41 41.22 BioMistral ROUGE-1 42.16 49.33 52.12 41.64 24.27 41.11 41.77 InternLM 2 ROUGE-1 44.12 51.58 53.46 42.00 24.05 41.96 42.86 Llama 3 ROUGE-1 43.66 49.42 53.98 42.49 25.07 42.44 42.84 MMedLM (Ours) ROUGE-1 41.98 49.28 52.34 41.41 24.67 41.35 41.84 MMedLM 2 (Ours) ROUGE-1 44.77 51.78 57.10 45.30 28.18 44.59 45.29 MMed-Llama 3 (Ours) ROUGE-1 42.47 49.36 55.06 42.85 26.67 43.32 43.29 BLOOMZ ROUGE-2 15.25 19.78 21.22 16.63 6.13 16.44 15.91 InternLM ROUGE-2 16.19 19.86 22.13 14.67 5.64 14.92 15.57 Llama 2 ROUGE-2 15.75 18.07 22.81 15.76 6.33 15.81 15.76 ChatDoctor ROUGE-2 16.70 17.67 20.65 16.16 7.12 16.64 15.82 MedAlpaca ROUGE-2 16.41 17.63 21.78 16.21 7.11 16.49 15.94 PMC-LLaMA ROUGE-2 17.40 17.56 21.00 16.20 6.77 16.47 15.90 Meditron ROUGE-2 17.00 19.42 23.68 17.26 7.20 16.91 16.91 Mistral ROUGE-2 15.64 18.71 22.64 16.64 7.31 16.71 16.27 BioMistral ROUGE-2 15.76 20.49 24.00 17.44 7.64 16.75 17.01 InternLM 2 ROUGE-2 18.04 21.75 25.32 17.76 7.02 17.36 17.88 Llama 3 ROUGE-2 17.75 20.18 25.50 18.13 7.78 18.32 17.94 MMedLM (Ours) ROUGE-2 16.31 20.00 24.50 17.82 7.57 17.56 17.29 MMedLM 2 (Ours) ROUGE-2 18.47 22.56 28.88 20.80 10.02 19.68 20.07 MMed-Llama 3 (Ours) ROUGE-2 17.21 20.52 26.75 18.36 8.77 18.77 18.40 BLOOMZ ROUGE-L 36.53 32.84 32.03 37.17 19.90 35.94 32.40 InternLM ROUGE-L 37.89 33.59 32.49 34.37 20.13 34.26 32.12 Llama 2 ROUGE-L 37.05 31.09 32.45 36.44 21.73 35.72 32.41 ChatDoctor ROUGE-L 37.91 31.04 31.62 36.29 22.02 36.38 32.54 MedAlpaca ROUGE-L 38.00 31.43 32.08 36.57 21.77 36.13 32.66 PMC-LLaMA ROUGE-L 38.77 30.85 31.42 36.57 21.13 36.19 32.49 Meditron ROUGE-L 38.47 33.60 33.94 37.23 22.59 36.51 33.72 Mistral ROUGE-L 37.22 32.06 31.78 36.88 23.00 36.35 32.88 BioMistral ROUGE-L 37.61 33.89 33.13 37.45 22.62 36.50 33.53 InternLM 2 ROUGE-L 39.61 36.27 35.06 38.17 22.74 37.24 34.85 Llama 3 ROUGE-L 39.22 33.86 35.37 38.22 23.42 37.75 34.64 MMedLM (Ours) ROUGE-L 37.84 34.35 34.51 37.81 23.19 37.08 34.13 MMedLM 2 (Ours) ROUGE-L 40.19 36.97 38.31 41.17 26.62 39.71 37.16 MMed-Llama 3 (Ours) ROUGE-L 38.09 34.37 36.54 38.59 24.77 38.49 35.14
@@
[b] Method English Chinese Japanese French Russian Spanish Avg. BLOOMZ 74.72 76.46 67.87 76.36 67.24 75.58 73.04 InternLM 75.13 76.95 67.22 72.46 67.87 72.94 72.09 Llama 2 76.50 76.38 74.81 77.77 74.30 76.54 76.05 ChatDoctor 76.53 75.23 72.31 77.34 74.25 76.66 75.39 MedAlpaca 76.30 75.43 71.33 77.36 74.14 76.43 75.16 PMC-LLaMA 76.84 75.17 71.28 77.15 73.77 76.35 75.09 Meditron 76.57 75.13 70.96 76.27 72.43 75.67 74.51 Mistral 77.99 77.46 74.11 78.67 75.07 77.77 76.85 BioMistral 77.23 76.22 71.53 77.63 71.80 77.03 75.24 Llama 3 77.36 77.96 75.12 78.26 73.66 77.71 76.68 InternLM 2 77.36 78.75 73.05 77.58 73.68 76.66 76.18 MMedLM (Ours) 75.68 77.16 72.15 76.41 73.87 75.42 75.12 MMedLM 2 (Ours) 77.59 78.42 75.79 78.47 76.07 77.65 77.33 MMed-Llama 3 (Ours) 77.42 78.02 76.09 78.41 75.04 78.00 77.16
@@
[b] Method English Chinese Japanese French Russian Spanish Avg. BLOOMZ 76.66 78.30 74.26 78.18 72.22 76.92 76.09 InternLM 77.11 78.43 74.73 77.04 71.91 76.04 75.88 Llama 2 77.49 77.78 75.99 78.45 74.60 77.21 76.92 ChatDoctor 77.47 77.18 72.65 78.16 74.69 77.19 76.22 MedAlpaca 77.53 77.31 74.39 78.23 74.85 77.24 76.59 PMC-LLaMA 77.98 77.22 74.15 78.15 73.94 76.96 76.40 Meditron 77.84 78.38 76.13 79.00 75.14 77.43 77.32 Mistral 77.79 78.49 76.00 79.00 75.77 77.80 77.47 BioMistral 77.83 78.97 76.33 79.00 75.35 77.75 77.54 Llama 3 78.33 78.69 77.06 79.51 75.45 78.05 77.85 InternLM 2 78.44 79.37 76.70 79.38 74.95 78.01 77.81 MMedLM (Ours) 77.58 78.66 76.02 78.71 75.42 77.36 77.29 MMedLM 2 (Ours) 78.74 79.56 78.40 80.36 77.41 78.73 78.87 MMed-Llama 3 (Ours) 78.05 78.96 77.75 79.76 75.94 78.30 78.13
@@
E.3 Detailed Results on Rating Scores
In Supplementary Table 13 and Supplementary Table 14, we further show the GPT-4 and human rating scores for six languages respectively.
[b] Method English Chinese Japanese French Russian Spanish Avg. MMed-Llama 3 (Ours) 4.14 4.32 4.24 4.38 3.64 3.86 4.10 InternLM 2 3.88 3.82 3.86 3.78 3.98 3.98 3.88 Llama 3 3.44 3.36 3.94 3.48 3.72 3.56 3.58 BioMistral 3.54 3.92 4.02 3.46 4.04 3.72 3.78 MEDITRON 3.60 3.18 2.84 3.16 3.50 3.40 3.28 GPT-3.5 2.40 2.40 2.10 2.74 2.12 2.46 2.37
[b] Method English Chinese Japanese French Russian Spanish Avg. MMed-Llama 3 (Ours) 4.24 5.18 4.80 4.72 4.56 4.86 4.73 InternLM 2 3.92 4.30 3.62 3.64 3.52 4.02 3.84 Llama 3 3.46 3.48 3.56 3.94 3.56 3.40 3.57 BioMistral 3.24 3.30 3.62 2.90 3.94 3.54 3.42 MEDITRON 3.66 2.58 2.48 3.02 3.02 2.72 2.91 GPT-3.5 2.48 2.16 2.92 2.78 2.40 2.46 2.53